# Probabilistic Approaches to The Energy Equality in Forced Surface Quasi-Geostrophic Equations

### v1

@@ 3. Existence and Exponential Tightness
$\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $=\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (3.8)
$+\varepsilon^{1/2}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle+\frac{\varepsilon}{2}t\left\|P_{m}\Lambda^{\alpha}\right\|_{HS}^{2}.$
It follows from (3.8), Burkholder-Davis-Gundy inequality and the scaling regime (2.8) that
$\mathbb{E}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\mathbb{E}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\leqslant\mathbb{E}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+CT.$ (3.9)
This implies that $\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle$ has bounded quadratic variation, and therefore it is a martingale. Consequently, for any $\eta>0$, the process
$\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}$
is a submartingale. By Doob’s maximal inequality, we see that
$\mathbb{E}\left[\sup_{t\in[0,T]}\right.$ $\left.\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.10)
$\leqslant 4\mathbb{E}\left[\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right].$
Taking the exponential function of both sides of (3.8) and using (2.8) yield that
$\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.11)
$\leqslant$ $\exp\big{\{}\frac{C(\eta)}{\varepsilon}T\big{\}}\cdot\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}.$
@@
$\quad\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big{\}}\Big{)}$ (3.12)
$=\sum_{0<|k|\leqslant m}\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{\varepsilon}|k|^{2\alpha-4\beta}\langle e_{k},\theta_{\varepsilon,m}(0)\rangle^{2}\Big{\}}\Big{)}$
$=\sum_{0<|k|\leqslant m}\varepsilon\log(1-\eta)^{-1/2}\leqslant\varepsilon\sum_{0<|k|\leqslant m}1\leqslant C(\eta).$
@@
Taking supremum and expectation to (3.11), combining with Hölder’s inequality, (3.10) and (3.12), we obtain
@@
$\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.13)
@@
$=$ $\mathbb{E}\left[\sup_{t\in[0,T]}\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]$
@@
$\leqslant$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\mathbb{E}\exp\bigg{[}\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}$
@@
$\cdot$ $\sup_{t\in[0,T]}\left(\exp\left\{-\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right)\bigg{]}$
@@
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{\frac{2\eta}{\varepsilon}\left\|\theta_{\varepsilon,m}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\right]^{1/2}$
@@
$\cdot$ $\left[\mathbb{E}\sup_{t\in[0,T]}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
@@
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
@@
By the covariance structure of Brownian motions,
$-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle$
has the quadratic variation process
@@
$\left\langle\left\langle-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{\cdot}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\rangle\right\rangle_{t}=\frac{16\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
@@
Thus $\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{8\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ is a local martingale. Moreover, for any $\eta\in(0,1/8)$,
@@
$M_{2}(t):=\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
@@
$\mathbb{E}\left[\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.14)
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\cdot\left[\mathbb{E}\exp\left\{-\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{t}\ \left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}.$
@@
$\quad\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
$\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{-\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m}(s),P_{m}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
$\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
@@
$\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
@@
$\varepsilon\log\mathbb{P}\left(\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\right)$
$\leqslant$ $-\varepsilon R+\varepsilon\log\mathbb{E}\left(\frac{\eta}{\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{2\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)$
$\leqslant$ $-\varepsilon R+C(\eta)T.$
@@
The proof of the exponential estimates for $\theta_{\varepsilon,\delta}$ can be handled in much the same way under the scaling regime (2.7) with initial data satisfying Assumption A. The only difference is that we need to replace the finite-dimensional Itô’s formula by the pathwise $H^{\alpha-2\beta}$-energy equality (3.1). In contrast, we prove the following exponential estimates for $\theta_{\varepsilon,\delta}$ under the scaling regime (2.7). These results hold for $\theta_{\varepsilon,m}$ under the scaling regime (2.8) by the same arguments.
@@
Let $\varepsilon,\delta>0$. Suppose that $\theta_{\varepsilon,\delta}$ is a stochastic generalized Leray solution of (1.3) in the sense of Definition 3.1, with initial data $\theta_{\varepsilon,\delta}(0)$ satisfying Assumption A. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (2.7), we have
$\varepsilon\log\mathbb{E}\Big{(}\exp\Big{\{}\frac{\eta}{2\varepsilon}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big{\}}\Big{)}\leqslant C(\eta,T),$ (3.15)
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big{(}\|\theta_{\varepsilon,\delta}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big{)}=-\infty.$ (3.16)

### v3

@@ 3. Existence and Exponential Tightness
$\frac{1}{2}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $=\frac{1}{2}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}-\int_{0}^{t}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (3.2)
$+\varepsilon^{1/2}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle+\frac{\varepsilon}{2}t\left\|P_{m(\varepsilon)}\Lambda^{\alpha}\right\|_{HS}^{2}.$
It follows from (3.2), Burkholder-Davis-Gundy inequality, and the scaling regime (2.5) that
$\frac{1}{2}\mathbb{E}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}$ $+\mathbb{E}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s$ (3.3)
$\leqslant\frac{1}{2}\mathbb{E}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+CT.$
For any $\eta>0$, the process $\exp\left\{\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}$ is a submartingale. By Doob’s maximal inequality, we see that
$\mathbb{E}\left[\sup_{t\in[0,T]}\right$. $\left.\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.4)
$\leqslant 4\mathbb{E}\left[\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right].$
For any $\varepsilon$ and $\eta>0$, taking the exponential function of both sides of (3.2) and using (2.5) yield that
$\exp\left\{\frac{\eta}{2\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.5)
$\leqslant$ $\exp\big\{\frac{C(\eta)}{\varepsilon}T\big\}\cdot\exp\left\{\frac{\eta}{2\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\cdot\exp\left\{\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}.$
@@
$\quad\varepsilon\log\mathbb{E}\Big(\exp\Big\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\Big\}\Big)$ $=\sum_{0<|k|\leqslant m(\varepsilon)}\varepsilon\log\mathbb{E}\Big(\exp\Big\{\frac{\eta}{\varepsilon}|k|^{2\alpha-4\beta}\langle e_{k},\theta_{\varepsilon,m(\varepsilon)}(0)\rangle^{2}\Big\}\Big)$ (3.6)
$=\sum_{0<|k|\leqslant m(\varepsilon)}\varepsilon\log(1-\eta)^{-1/2}\leqslant\varepsilon\sum_{0<|k|\leqslant m(\varepsilon)}1\leqslant C(\eta).$
@@
Taking supremum and expectation to (3.5), combining with Hölder’s inequality, (3.4), and (3.6), we obtain
@@
$\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ (3.7)
@@
$=$ $\mathbb{E}\left[\sup_{t\in[0,T]}\exp\left\{\frac{\eta}{2\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]$
@@
$\leqslant$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\mathbb{E}\exp\bigg[\left\{\frac{\eta}{2\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}$
@@
$\cdot$ $\sup_{t\in[0,T]}\left(\exp\left\{\frac{\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right)\bigg]$
@@
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\left\|\theta_{\varepsilon,m(\varepsilon)}(0)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}\right\}\right]^{1/2}$
@@
$\cdot$ $\left[\mathbb{E}\sup_{t\in[0,T]}\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
@@
$\lesssim$ $\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}$
@@
By the covariance structure of Brownian motions and the definition of the quadratic variation process, it holds that
@@
$\left\langle\left\langle\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{\cdot}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\rangle\right\rangle_{t}=\frac{16\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s.$
@@
Thus $\exp\left\{\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle-\frac{8\eta^{2}}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$ is a local martingale. Moreover, for any $\eta\in(0,1/8)$,
@@
$M_{1}(t):=\exp\left\{\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{t}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{t}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
@@
$\mathbb{E}\left[\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right]$ (3.8)
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\cdot\left[\mathbb{E}\exp\left\{\frac{4\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle-\frac{\eta}{\varepsilon}\int_{0}^{T}\ \left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}$
$\lesssim$ $\left[\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\right]^{1/2}.$
@@
$\quad\mathbb{E}\exp\left\{\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}$
$\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}\left[\mathbb{E}\exp\left\{\frac{2\eta}{\varepsilon^{1/2}}\int_{0}^{T}\left\langle\Lambda^{2\alpha-2\beta}\theta_{\varepsilon,m(\varepsilon)}(s),P_{m(\varepsilon)}\mathrm{d}W(s)\right\rangle\right\}\right]^{1/2}\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
@@
$\mathbb{E}\exp\left\{\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right\}\leqslant\exp\left\{\frac{C(\eta)}{\varepsilon}T\right\}.$
@@
$\varepsilon\log\mathbb{P}\left(\frac{1}{2}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\geqslant R\right)$
$\leqslant$ $-\eta R+\varepsilon\log\mathbb{E}\left(\frac{\eta}{2\varepsilon}\sup_{t\in[0,T]}\left\|\theta_{\varepsilon,m(\varepsilon)}(t)\right\|_{H^{\alpha-2\beta}(\mathbb{T}^{2})}^{2}+\frac{\eta}{\varepsilon}\int_{0}^{T}\left\|\theta_{\varepsilon,m(\varepsilon)}(s)\right\|_{H^{2\alpha-2\beta}(\mathbb{T}^{2})}^{2}\,\mathrm{d}s\right)$
$\leqslant$ $-\eta R+C(\eta)T.$
@@
The proof of the exponential estimates for $\theta_{\varepsilon,\delta(\varepsilon)}$ can be handled in much the same way under the scaling regime (1.7) with initial data satisfying Assumption A. The only difference is that we need to replace the finite-dimensional Itô’s formula by the pathwise $H^{\alpha-2\beta}$-energy equality (1.9). In contrast, we prove the following exponential estimates for $\theta_{\varepsilon,\delta(\varepsilon)}$ under the scaling regime (1.7). These results hold for $\theta_{\varepsilon,m(\varepsilon)}$ under the scaling regime (2.5) by the same arguments.
@@
Let $\varepsilon,\delta(\varepsilon)>0$. Suppose that $\theta_{\varepsilon,\delta(\varepsilon)}$ is a stochastic generalized Leray solution of (1.3) in the sense of Definition 1.2, with initial data $\theta_{\varepsilon,\delta(\varepsilon)}(0)$ satisfying Assumption A. Fix $\sigma\in(0,1/2)$ and $l>3+2\beta-\alpha$. Then there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0}]$, under the scaling regime (1.7), we have
$\varepsilon\log\mathbb{E}\Big(\exp\Big\{\frac{\eta}{\varepsilon}\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}^{2}\Big\}\Big)\leqslant C(\eta,T),$ (3.9)
@@
$\limsup_{R\rightarrow\infty}\sup_{\varepsilon\in(0,1)}\varepsilon\log\mathbb{P}\Big(\|\theta_{\varepsilon,\delta(\varepsilon)}\|_{W^{\sigma,2}([0,T];H^{-l}(\mathbb{T}^{2}))}>R\Big)=-\infty.$ (3.10)